An airplane flying with a tailwind takes 8 hours to fly a distance of 1,600miles. If the return trip, against the wind takes 10 hours, find the speed of the plane and wind.
Answer by rfer(16322)

plane travels at 180mi/hr
wind is 20mi/hr
1600/10=160
1600/8=200
the difference of 40/2= wind of 20
160+20=180
200-20=180